# Optimal Stopping with a Predicted Prior

### v1

@@ 3.3 Brief Discussion on Algorithms Outside the Framework
In terms of MaxExp, the literature has considered algorithms that accept the first value greater than a threshold at its arrival time, regardless of whether the value is best-so-far. With a correct prior, the optimal policy by backward induction takes this form, where the threshold is the expected accepted value if the algorithm rejects the current value. For example, see the work by krengel1977semiamarts on general prophet inequality and the more recent research [correaPostedPriceMechanisms2017, liuVariableDecompositionProphet2021, harbNewProphetInequalities2025] for i.i.d. prophet inequality. These algorithms are outside our framework, where the taken value must be best-so-far, in addition to surpassing the time-dependent threshold.
We stress that the best-so-far requirement is critical for robustness. The mentioned algorithms that have no best-so-far requirements are only 0-robust. To see this, consider a predicted prior supported on values much smaller than those in the real prior. Then, all values from the real prior will have cdf value essentially $1$ w.r.t. the predicted prior. As a result, the algorithm accepts the first value arrived when $\theta(t)<1$, which is noncompetitive and leads to 0 robustness.
Meanwhile, our framework is limited in that it does not adjust thresholds adaptively based on the observed values. Intuitively, a “learning” algorithm may adjust thresholds toward Dynkin’s algorithm [dynkinOptimumChoiceInstant1963] if the observed values are unlikely to be from the predicted prior, and conversely, toward the $0.5801$-competitive algorithm [gilbertRecognizingMaximumSequence] in the MaxProb setting or the $0.745$-competitive algorithm [correaProphetInequalitiesIID2019] in the MaxExp setting if the observed values are consistent with the predicted prior. We consider a static threshold function to make the analyses tractable. Exploring algorithms with adaptive thresholds is an interesting direction for future work.
@@ 4 Implicit Sharding
For any $n,k\in\mathbb{N}$, and for both the MaxExp and MaxProb objectives, if there is a threshold function $\theta$ such that algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ is $\alpha$-consistent and $\beta$-robust for $nk$ values, then there is a corresponding $\alpha$-consistent $\beta$-robust algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values.
By our convention, $\tilde{F}^{\nicefrac{{1}}{{k}}}$ denotes the distribution defined by cdf $\tilde{F}^{\nicefrac{{1}}{{k}}}(\cdot)$. Importantly, the maximum of $k$ values drawn i.i.d. from $\tilde{F}^{\nicefrac{{1}}{{k}}}$ follows distribution $\tilde{F}$. We will construct an algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values to mimic the decision of algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ with $nk$ values drawn i.i.d. from $\tilde{F}^{\nicefrac{{1}}{{k}}}$. Let $x_{1},\dots,x_{n}$ be the $n$ values by their arrival order in the original instance. Further, consider an instance of $nk$ values $\{x^{\prime}_{ij}\}_{i\in[n],j\in[k]}$ drawn i.i.d. from distribution $F^{\nicefrac{{1}}{{k}}}$ with sorted arrival times $t^{\prime}_{11}<\cdots<t^{\prime}_{1k}<\cdots<t^{\prime}_{n1}<\cdots<t^{\prime}_{nk}$ drawn i.i.d. and uniformly from $[0,1]$.
Definition 6 (Coupling for Sharding).
Let $x_{i}=\max_{j\in[k]}x^{\prime}_{ij}$, and let $s_{i}$ be the arrival time of $\operatorname*{arg\,max}_{j\in[k]}x^{\prime}_{ij}$. We use $s_{i}$ instead of $t_{i}$ to avoid confusion, as $s_{i}$ is not uniform over $[0,1]$.
Naïvely, we would like to simulate algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ by sampling $x^{\prime}_{ij}$, $j\in[k]$, from $F^{\nicefrac{{1}}{{k}}}$ conditioned on the realization of $x_{i}$, and accept value $x_{i}$ if and only if one of the $x^{\prime}_{ij}$’s is accepted by $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$. This is the sharding technique by harbNewProphetInequalities2025 for the prophet inequality setting. Unfortunately, the algorithm does not know $F$ in the problem of optimal stopping with a predicted prior, and thus cannot sample $x^{\prime}_{ij}$ from $F^{\nicefrac{{1}}{{k}}}$. Sampling $x^{\prime}_{ij}$ based on the predicted prior $\tilde{F}$ and the corresponding $\tilde{F}^{\nicefrac{{1}}{{k}}}$ preserves the consistency of $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$, but not robustness.
We propose an implicit sharding method that partially simulates algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ only at the arrival times of the $x_{i}=\max_{j\in[k]}x^{\prime}_{ij}$. We will show that this partial simulation is at least as good as the full simulation in the original sharding method. Further, the partial simulation does not rely on $F$ and will not explicitly construct the $x^{\prime}_{ij}$’s, as they are only used in the analysis.
@@
2. For each $i\in[n]$, sample $s_{i}$ uniformly from $\{t^{\prime}_{i1},t^{\prime}_{i2},\ldots,t^{\prime}_{ik}\}$. 3. For each value $x_{i}$ when no value has been accepted, accept $x_{i}$ if: (a) Value $x_{i}$ is best-so-far; and (b) The cdf value of $x_{i}$ w.r.t. $\tilde{F}^{\nicefrac{{1}}{{k}}}$ is greater than the threshold at time $s_{i}$, i.e., if:
@@
It suffices to show that for any real prior $F$, and any realization of the values and arrival times under the defined coupling, the value accepted by algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ is greater than or equal to the value accepted by algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$.
Suppose algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ accepts value $x_{i}$, and $x^{\prime}_{ij}=x_{i}$ is the corresponding value in the coupled instance for algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$. Since $x_{i}$ is best-so-far, $x^{\prime}_{ij}$ is also best-so-far because
$x^{\prime}_{ij}=x_{i}\geq x^{\prime}_{i^{\prime}j^{\prime}}\quad\text{for any }i^{\prime}\leq i,\;j^{\prime}\in[k].$
@@
Hence, algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ accepts either $x^{\prime}_{ij}=x_{i}$, or a value that arrives earlier, which must be at most $x^{\prime}_{ij}=x_{i}$ since $x_{i}$ is best-so-far. Taking expectations over the coupling completes the proof. ∎
@@ 5 Proof of Robustness
We analyze the robustness of our algorithm for MaxProb (which then implies robustness for MaxExp) in a stronger setting that fixes the realized values and randomizes only the arrivals. This problem is also known as the game of Googol [fergusonWhoSolvedSecretary1989, gnedin1994solution].
@@
For any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1},x_{2},\dots,x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value with probability at least
@@
This implies the robustness guarantees for both MaxProb and MaxExp as corollaries through a standard argument.
Corollary 9.
For any predicted prior $\tilde{F}$, any threshold function $\theta$, and any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ is $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$-robust for both MaxProb and MaxExp.
Proof.
@@
$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~=~\operatorname{\mathbf{E}}_{x_{1},\dots,x_{n}}\Big[\,\operatorname{\mathbf{Pr}}_{t_{1},\dots,t_{n}}\big[\,x_{\textsc{Alg}}=x_{\textsc{Opt}}\mid x_{1},\dots,x_{n}\,\big]\,\Big]~.$
By Proposition 8, the inner probability is at least $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$ for every realization, so the expectation is at least the same quantity. This proves robustness for MaxProb. Moreover, it implies $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]\geq\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}\cdot\operatorname{\mathbf{E}}[x_{\textsc{Opt}}]$, i.e., the same robustness for MaxExp. ∎
@@
For any $n\in\mathbb{N}$, any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1}<x_{2}<\dots<x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value $x_{n}$ with probability
@@
Since $x_{n}$ is always best-so-far, the algorithm accepts $x_{n}$ if and only if 1) its cdf value under $\tilde{F}$ is greater than the threshold, i.e., $\tilde{F}(x_{n})>\theta(t_{n})$, and 2) the algorithm accepts no value before $t_{n}$. By Proposition 4, the latter is equivalent to the event that the prefix maximum for time $t_{n}$ has cdf under $\tilde{F}$ at most the threshold at its arrival time. Hence, the acceptance probability is
@@
To compute the expression for the probability in the integrand, consider each possible prefix maximum separately. First, the prefix maximum is 0 if no value arrives before $t_{n}$, which occurs with probability $(1-t_{n})^{n-1}$. Further, the prefix maximum is $x_{i}$ if $x_{i}$ arrives at time $t_{i}\in[0,t_{n})$, and the larger values $x_{i+1},\ldots,x_{n-1}$ arrive after time $t_{n}$, which happens with probability $(1-t_{n})^{n-1-i}$. In this case, we also require $\tilde{F}(x_{i})\leq\theta(t_{i})$ for the arrival time $t_{i}\in[0,t_{n})$, contributing
@@
Putting together and changing variables $t_{n}$ with $t$ and $t_{i}$ with $s$, the lemma follows. ∎
@@
First, we fix $\theta$ and characterize the worst-case predicted prior $\tilde{F}$. Equation 2 depends on $\tilde{F}$ through $n$ values $0\leq\tilde{F}(x_{1})\leq\dots\leq\tilde{F}(x_{n})\leq 1$. Further, it is non-decreasing in $\tilde{F}(x_{n})$, and non-increasing in $\tilde{F}(x_{1}),\dots,\tilde{F}(x_{n-1})$. Therefore, given any fixed threshold function $\theta$, Equation 2 is minimized when all $\tilde{F}(x_{i})$ are equal to some $q\in[0,1]$.
In this case, the threshold function is effectively binary. A threshold $\theta(t)<q$ is equivalent to $\theta(t)=0$, as $\tilde{F}(x_{1}),\dots,\tilde{F}(x_{n})$ all surpass the threshold. Similarly, a threshold $\theta(t)\geq q$ is equivalent to $\theta(t)=1$, as none surpasses it. Hence, we may without loss of generality consider a binary threshold, such that for some $\lambda\in[\lambda_{1},\lambda_{2}]$, it holds
@@
This corresponds to Dynkin’s algorithm with a suboptimal transition point of $\lambda$ instead of the optimal $\frac{1}{\mathrm{e}}$. This is $(-\lambda\ln\lambda)$-competitive (e.g., feldmanImprovedCompetitiveRatios2011). Since $\lambda\ln\lambda$ is concave in $\lambda$ on $[0,1]$, its minimum over $\lambda_{1},\lambda_{2}]$ is attained at an endpoint, yielding the bound $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$. ∎

### v2

@@ 3.4 Brief Discussion on Algorithms Outside the Framework
In terms of MaxExp, the literature has considered algorithms that accept the first value greater than a threshold at its arrival time, regardless of whether the value is a record. With a correct prior, the optimal policy by backward induction takes this form, where the threshold is the expected accepted value if the algorithm rejects the current value. For example, see the work by Krengel and Sucheston [42] on general prophet inequality and the more recent research [16, 44, 37] for i.i.d. prophet inequality. These algorithms are outside our framework, where the taken value must be a record, in addition to surpassing the time-dependent threshold.
We stress that record-based requirement is critical for robustness. The mentioned algorithms without record-based requirements are only 0-robust. To see this, consider a predicted prior supported on values much smaller than those in the real prior. Then, all values from the real prior will have a cdf value essentially $1$ w.r.t. the predicted prior. As a result, the algorithm accepts the first value that arrives when $\theta(t)<1$, which is noncompetitive and leads to 0 robustness.
Meanwhile, our framework is limited in that it does not adjust thresholds adaptively based on the observed values. Intuitively, a “learning” algorithm may adjust thresholds toward Dynkin’s algorithm [21] if the observed values are unlikely to be from the predicted prior, and conversely, toward the $0.580$-competitive algorithm [33] in the MaxProb setting or the $0.745$-competitive algorithm [14] in the MaxExp setting if the observed values are consistent with the predicted prior. We consider a static threshold function to make the analyses tractable. Exploring algorithms with adaptive thresholds is an interesting direction for future work.
@@ 4 Implicit Sharding
For any $n,k\in\mathbb{N}$, for any predicted prior $\tilde{F}$, and for both the MaxExp and MaxProb objectives, if there is a threshold function $\theta$ such that algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ is $\alpha$-consistent and $\beta$-robust for $nk$ values, then there is a corresponding $\alpha$-consistent $\beta$-robust algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values.
By convention, $\tilde{F}^{\nicefrac{{1}}{{k}}}$ denotes the distribution defined by cdf $\tilde{F}^{\nicefrac{{1}}{{k}}}(\cdot)$. Importantly, the maximum of $k$ values drawn i.i.d. from $\tilde{F}^{\nicefrac{{1}}{{k}}}$ follows distribution $\tilde{F}$. We will construct an algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values $x_{1},x_{2},\ldots,x_{n}$ to mimic the decision of algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ with $nk$ values $x^{\prime}_{11},\ldots,x^{\prime}_{1k},\ldots,x^{\prime}_{n1},\ldots,x^{\prime}_{nk}$ drawn i.i.d. from $\tilde{F}^{\nicefrac{{1}}{{k}}}$.
Naïvely, we would like to simulate algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ by sampling $x^{\prime}_{ij}$, $j\in[k]$, from $F^{\nicefrac{{1}}{{k}}}$ conditioned on the realization of $x_{i}$, and accept value $x_{i}$ if and only if one of the $x^{\prime}_{ij}$’s is accepted by $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$. This is the sharding technique by Harb [37] for the prophet inequality setting. Unfortunately, the algorithm does not know $F$ in the problem of optimal stopping with a predicted prior, and thus cannot sample $x^{\prime}_{ij}$ from $F^{\nicefrac{{1}}{{k}}}$. Sampling $x^{\prime}_{ij}$ based on the predicted prior $\tilde{F}$ and the corresponding $\tilde{F}^{\nicefrac{{1}}{{k}}}$ preserves the consistency of $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$, but not robustness.
We propose an implicit sharding method that partially simulates algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ only at the arrival times of the $x_{i}=\max_{j\in[k]}x^{\prime}_{ij}$. We will show that this partial simulation is at least as good as the full simulation in the original sharding method. In contrast to the sharding technique, the partial simulation does not rely on $F$ and will not explicitly construct the $x^{\prime}_{ij}$’s, as they are only used in the analysis. Our algorithm is defined as follows.
@@
2. For each $i\in[n]$, sample $s_{i}$ uniformly from $\{t^{\prime}_{i1},t^{\prime}_{i2},\ldots,t^{\prime}_{ik}\}$. 3. For each value $x_{i}$ when no value has been accepted, accept $x_{i}$ if (a) Value $x_{i}$ is a record; and (b) The cdf value of $x_{i}$ w.r.t. $\tilde{F}^{\nicefrac{{1}}{{k}}}$ is greater than the threshold at time $s_{i}$, i.e., if
@@
Next, we design a coupling that gives a joint distribution preserving the marginal distributions of the random variables used in our algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ and in the simulated algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$, respectively.
Definition 6 (Coupling).
Sample $nk$ i.i.d. values from $F^{\nicefrac{{1}}{{k}}}$. For each value, sample an arrival time i.i.d. from $\mathrm{Unif}[0,1]$. Then, we index the values by the ascending order of their arrival times. Denote the values by $\{x^{\prime}_{ij}\}_{i\in[n],j\in[k]}$ and the sorted times by
For each $i$, define
$x_{i}\coloneqq\max_{j\in[k]}x^{\prime}_{ij},\qquad s_{i}\coloneqq t^{\prime}_{ij_{i}},\quad\text{where }j_{i}\in\operatorname*{arg\,max}_{j\in[k]}x^{\prime}_{ij}.$
Under this coupling, each $x_{i}$ has distribution $F$, and $\max_{i\in[n]}x_{i}=\max_{i\in[n],j\in[k]}x^{\prime}_{ij}$ pointwise. Also, because the shards in each block are i.i.d., the argmax index $j_{i}$ is uniform over $[k]$, so $s_{i}$ has the same distribution as Step 2 in $\mathcal{A}_{\mathrm{sharding}}$.
We stress that the algorithm does not rely on $F$ and will not explicitly construct the $x^{\prime}_{ij}$’s. They are only used in the coupling for analysis.
@@
We establish the pointwise value dominance property: for any prior $F$, and any realization of the values and arrival times under the defined coupling, the value accepted by algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ is greater than or equal to the value accepted by algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$.
Suppose algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ accepts value $x_{i}$, and $x^{\prime}_{ij}=x_{i}$ is the corresponding value in the coupled instance for algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$. Since $x_{i}$ is a record, $x^{\prime}_{ij}$ is also a record because
$x^{\prime}_{ij}=x_{i}\geq x^{\prime}_{i^{\prime}j^{\prime}},t^{\prime}_{ij}=s_{i}\geq t^{\prime}_{i^{\prime}j^{\prime}}\quad\text{for any }i^{\prime}<i,\;j^{\prime}\in[k]\text{ or }i^{\prime}=i,\;j^{\prime}\leq j.$
@@
Hence, algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ accepts either $x^{\prime}_{ij}=x_{i}$, or a value that arrives earlier, which must be at most $x^{\prime}_{ij}=x_{i}$ since $x_{i}$ is a record.
By this pointwise value dominance, we obtain the following conclusion for each objective:
  1. 1.
    For MaxExp: The conclusion follows directly by taking expectations over the coupling.
  2. 2.
    For MaxProb: The conclusion follows directly by taking expectations over all points in the coupling where $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ accepts the maximum value.
This completes the proof. ∎
@@ 5 Proof of Robustness
In Section 3 we stated Lemma 7, which bounds the robustness of $\mathcal{A}(\tilde{F},\theta)$ in our stochastic model. We prove it in this section. The argument proceeds via a stronger “fixed values, random arrivals” model, also known as the game of Googol [29, 34], in which the values $x_{1},x_{2},\ldots,x_{n}$ are treated as deterministic and only the arrival order is random.
The key step is the following bound in the game-of-Googol setting.
@@
For any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1},x_{2},\dots,x_{n}$, algorithm $\mathcal{A}(\tilde{F},\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ accepts the maximum value with probability at least
@@
We now derive Lemma 7 from Proposition 10.
Proof of Lemma 7.
Fix a predicted prior $\tilde{F}$ and a threshold function $\theta$. By definition, for $\lambda_{1}=\sup\{0\leq t\leq 1:\theta(t)=1\}$ and $\lambda_{2}=\inf\{0\leq t\leq 1:\theta(t)=0\}$, we have $\theta=\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$. Thus, it suffices to show that $\mathcal{A}(\tilde{F},\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ is $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$-robust.
@@
$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~=~\operatorname{\mathbf{E}}_{x_{1},x_{2},\ldots,x_{n}}\big[\operatorname{\mathbf{Pr}}_{t_{1},t_{2},\ldots,t_{n}}\big[\,x_{\textsc{Alg}}=x_{\textsc{Opt}}\mid x_{1},x_{2},\ldots,x_{n}\,\big]\big]~.$
Applying Proposition 10, the inner probability is at least $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$ for every realization, so the expectation is at least the same quantity. This proves robustness for MaxProb. Moreover, it implies $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]\geq\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}\cdot\operatorname{\mathbf{E}}[x_{\textsc{Opt}}]$, since for any realization of the values, the accepted value equals the maximum value with the same probability as in the MaxProb case. This proves robustness for MaxExp as well. ∎
@@
For any $n\in\mathbb{N}$, any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1}<x_{2}<\dots<x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value $x_{n}$ with probability
@@
Since $x_{n}$ is always a record, the algorithm accepts $x_{n}$ if and only if 1) its cdf value under $\tilde{F}$ is greater than the threshold, i.e., $\tilde{F}(x_{n})>\theta(t_{n})$, and 2) the algorithm accepts no value before $t_{n}$. By Proposition 4, the latter is equivalent to the event that the prefix maximum for time $t_{n}$ has cdf under $\tilde{F}$ at most the threshold at its arrival time. Hence, the acceptance probability is
@@
To compute the expression for the probability in the integrand, consider each possible prefix maximum separately. First, the prefix maximum is 0 if no value arrives before $t_{n}$, which occurs with probability $(1-t_{n})^{n-1}$. Further, the prefix maximum is $x_{i}$ if $x_{i}$ arrives at time $t_{i}\in[0,t_{n})$, and the larger values $x_{i+1},x_{i+2},\ldots,x_{n-1}$ arrive after time $t_{n}$, which happens with probability $(1-t_{n})^{n-1-i}$. In this case, we also require $\tilde{F}(x_{i})\leq\theta(t_{i})$ for the arrival time $t_{i}\in[0,t_{n})$, contributing
@@
Putting everything together and changing variables $t_{n}$ to $t$ and $t_{i}$ to $s$, the lemma follows. ∎
@@
First, we fix $\theta$ and characterize the worst-case predicted prior $\tilde{F}$. Equation 3 depends on $\tilde{F}$ through $n$ values $0\leq\tilde{F}(x_{1})\leq\tilde{F}(x_{2})\leq\cdots\leq\tilde{F}(x_{n})\leq 1$. Further, it is non-decreasing in $\tilde{F}(x_{n})$, and non-increasing in $\tilde{F}(x_{1}),\tilde{F}(x_{2}),\ldots,\tilde{F}(x_{n-1})$. Therefore, given any fixed threshold function $\theta$, Equation 3 is minimized when all $\tilde{F}(x_{i})$ are equal to some $q\in[0,1]$.
In this case, the threshold function is effectively binary. A threshold $\theta(t)<q$ is equivalent to $\theta(t)=0$, as $\tilde{F}(x_{1}),\tilde{F}(x_{2}),\ldots,\tilde{F}(x_{n})$ all surpass the threshold. Similarly, a threshold $\theta(t)\geq q$ is equivalent to $\theta(t)=1$, as none surpasses it. Hence, we may without loss of generality consider a binary threshold, such that for some $\lambda\in[\lambda_{1},\lambda_{2}]$, it holds
@@
This corresponds to Dynkin’s algorithm with a suboptimal transition point of $\lambda$ instead of the optimal $\nicefrac{{1}}{{\mathrm{e}}}$. This algorithm is $(-\lambda\ln\lambda)$-competitive (e.g., Feldman et al. [27]). Since $\lambda\ln\lambda$ is concave in $\lambda$ on $[0,1]$, its minimum over $\lambda_{1},\lambda_{2}]$ is attained at an endpoint, yielding the bound $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$. ∎